OpenStudy (anonymous):

determine whether the graphs of the given equations are parallel, perpendicular, or neither. Explain. y=x+11 y=-2x+3

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

how do i determine whether they ar parallel or perpendicular?

OpenStudy (anonymous):

graph it?

OpenStudy (anonymous):

look only at the slopes of the given equations. If their slopes are similar, that means they are parallel. If their slopes are negative reciprocal of each other, that means they are perpendicular. If their slopes are neither of the above mentioned conditions, then they are neither parallel or perpendicular.

OpenStudy (anonymous):

*nor

OpenStudy (catlover5925):

perpendicular means crossing, or intersecting right??

OpenStudy (anonymous):

ok so the first one is y=x+11 and y=-x+2 this one is perpendicular correct?

OpenStudy (anonymous):

yes perpendicular is like this 90° to each other

OpenStudy (anonymous):

yes, that's right they are perpendicular because slope of the first line is 1 while the second one is -1

OpenStudy (catlover5925):

this is the first one graphed, by graphing you can tell if it is parallel or perpendicular

OpenStudy (anonymous):

y=-2x+3 2x+y=7 this is the second problem the slopes are -2 and ?

OpenStudy (anonymous):

this is the question posted above: let's check whether it is right..... y=x+11 -> slope is \(\sf m= 1\) y=-2x+3 -> slope is \(\sf m= -2\) therefore this is \(\sf neither\) parallel or perpendicular. @Catlover5925 you must also consider the angle between the lines, they should be 90 degrees to each other...

OpenStudy (anonymous):

i am so sorry i put the wrong equations next to each other

OpenStudy (anonymous):

the one i just put is the second problem

OpenStudy (anonymous):

@Khalimilan change the second equation to the form \(y=mx+b\) then you can easily determine what is the slope of the second line (:

OpenStudy (anonymous):

just rearrange the equation by subtracting both sides by 2x

OpenStudy (anonymous):

y=7-2

OpenStudy (anonymous):

y=7-2x or y=-2x+7, which means that the slope is ...?

OpenStudy (anonymous):

7 or 2x but isn't it 7? so would the answer be it is neither parallel or perpendicular?

OpenStudy (anonymous):

slope always goes with the variable 'x'

OpenStudy (anonymous):

so its 2x. so its perpendicular?

OpenStudy (anonymous):

you forget the negative (:

OpenStudy (anonymous):

oops :) but is the answer perpendicular since its 2x and -2x?

OpenStudy (anonymous):

let's see.. you have: y=-2x+3 2x+y=7 -> y=-2x+7 slope of the first equation is -2, and the second equation is -2, they look the same, so they should be?

OpenStudy (anonymous):

parallel!

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

you are a life saver haha thank you :)

OpenStudy (anonymous):

no problem (: do you understand how to do it now?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

remember the first problem you helped me with? i have two more problems that are pretty much the same and i still need help going step by step through them.

OpenStudy (anonymous):

btw, IF they are perpendicular, one slope should be -2 while the other slope should be \(\frac{1}{2}\) not 2, because we are looking the ''negative reciprocal" in perpendicular lines. okay, just tell me what you have so far, and i'll look through them (:

OpenStudy (anonymous):

thanks! do you want me to open up another question?